OpenStudy (anonymous):

to avoid a negative square root, how can i use the square root property on this ? (x+81)^2 = -158 or am i missing something about the rules of negative numbers and square roots ? I thought negative square roots were not allowed. thanks :)

OpenStudy (tkhunny):

If we allow only Real Numbers, we stop in our tracks and say, "No Solution".

OpenStudy (anonymous):

in math if it is a negative number we call it the number I because it is imaginary which is why we write something like this sqrt -158i if not allowed than no solution..

OpenStudy (anonymous):

Negative numbers result in imaginary numbers. Imaginary numbers scare me.

OpenStudy (anonymous):

it is nothing to be scared of

OpenStudy (anonymous):

the answer to the original equation (generated at math.com) is x= 11 and x=7. the original equation is : x^2 -18x +77 =0 my attempt at the quadratic formula produced : x=(18 ± 2√(61) / 2) i decided i had done something incorrectly and started completing the square from the original equation. can adding 158 to both sides be done while completing the square ? what happened to the reply about adding 158 ?

OpenStudy (anonymous):

okay I know what is wrong

OpenStudy (anonymous):

have you tried this A^2+B^2+C^2 this is the other form of qurad.

OpenStudy (anonymous):

it should be =C^2 not plus srry..

OpenStudy (anonymous):

trying it now.

OpenStudy (anonymous):

what you get for answer?

OpenStudy (anonymous):

1585=5929 i thought that formula looked familiar, so i searched for it and found he "Pythagorean theorem" http://en.wikipedia.org/wiki/Pythagorean_theorem perhaps, i have misunderstood something.

OpenStudy (anonymous):

@gemstonegirl come and see this

OpenStudy (anonymous):

need to know something does the problem state that no negative numbers can be the answer?

OpenStudy (anonymous):

ok, before the beginning of march, most of algebra was unknown to me. i only had the basic concept of replacing numbers with variables and solving for unknowns. this was required for calculations in ohms law when i was in collage. P=I*E for example. there was nothing complected about it , yet i struggled due to my lack of mathematical knowledge. i used an advanced TI calculator to solve what i did not comprehend. i dropped out of collage (foolish) because it all seemed overwhelming and i moved out of state with a girl that turned out to be insane. that was 5~ ago. when i came back , i found collage debt and a new opportunity with an apprenticeship program by IBEW (international brotherhood of electrical workers). there is an aptitude test near the end of April that involves algebra. i have been cramming my head with algebra ever since the day i learned of this. so, do not fret over your lack of algebra knowledge or mistakes, because i have less ten two months of algebra study ;)

OpenStudy (anonymous):

@ibew try this http://math.about.com/od/Equations/ss/Quadratic

OpenStudy (anonymous):

@ibew I amm studing math myself and this Summer I start on-line at IU East so math is my life and mistakes like this I should not make..

OpenStudy (anonymous):

@ibew there is no answer for this problem. also use the sqrt to see there is no answer to this problem ...

OpenStudy (anonymous):

is that allowed while "completing the square" ?

OpenStudy (anonymous):

the original equation is : x^2 -18x +77 =0

OpenStudy (anonymous):

i did: "ibew Medals 0 the answer to the original equation (generated at math.com) is x= 11 and x=7. the original equation is : x^2 -18x +77 =0 my attempt at the quadratic formula produced : x=(18 ± 2√(61) / 2) i decided i had done something incorrectly and started completing the square from the original equation. can adding 158 to both sides be done while completing the square ? what happened to the reply about adding 158 ? an hour ago" "

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@ibew there is no answer..

OpenStudy (anonymous):

this web site says there is an answer. http://www.algebrahelp.com/calculators/equation/calc.do;jsessionid=52F7E4DBE8C8F1538B2EE5187783E04C?equation=x^2+-18x+%2B77+%3D0+&solvf=AUTO i am looking it over now. i also added your link to my "math" book marks folder, thanks :)

OpenStudy (anonymous):

Your welcome anytime

OpenStudy (anonymous):

it does not use the quadratic formula or "completing the square" . i am nor well versed on factoring polynomials . i have memorizing the quadratic formula and completing the square because they achieve results 100% of the various problems.

OpenStudy (anonymous):

if any one can show me what i am doing wrong or how to get beyond this "soon to be" negative square root", it would be greatly appreciated. :)

OpenStudy (anonymous):

if any one has successfully solved the equation with the quadratic formula, that would help also.

OpenStudy (anonymous):

i think the decision is unanimous that there is no solution to this :D ure online answers may have been wrong ...never know

OpenStudy (anonymous):

@Luis_Rivera , thanks, i will post the question again with more information.

OpenStudy (anonymous):

lol sorry i dont know that far ..but u will have to change up the equation ...